OpenStudy (anonymous):

(7(-3)-3)/((-3)-3) Simplify the answer and please show the steps used to achieve it.

OpenStudy (lgbasallote):

first... what is 7(-3)?

OpenStudy (anonymous):

7 x (-3)

OpenStudy (lgbasallote):

right...and that is...?

OpenStudy (lgbasallote):

simplify 7(-3)

OpenStudy (anonymous):

-21

OpenStudy (anonymous):

and -21-3=-24 yes?

OpenStudy (lgbasallote):

yes

OpenStudy (lgbasallote):

now.. what is (-3) - 3?

OpenStudy (anonymous):

and for the bottom half of this (-3)-3=0?

OpenStudy (lgbasallote):

nope

OpenStudy (lgbasallote):

if -21 - 3 is -24...what is -3 -3?

OpenStudy (anonymous):

OH! now i get it!!! -3 -3=-6

OpenStudy (lgbasallote):

yes

OpenStudy (lgbasallote):

so you have -24/-6

OpenStudy (lgbasallote):

simplify...

OpenStudy (anonymous):

+4

OpenStudy (lgbasallote):

right

OpenStudy (lgbasallote):

so there's your answer

OpenStudy (anonymous):

Thank you so much for your help, and guiding me to answer.

OpenStudy (lgbasallote):

welcome